Prev Next

SW&IT Meeting Minutes 2022-08-08

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • management, catching up
    • irc centos6 rpm
    • jenkins upgrade breaks cemtos, rollback with DE/deam, docker centos6, swg-system PR
    • statserv/alh in dms
  • This week
    • PB lab checks with Brandon when back
    • statserv/alh in dms ctd.
    • dms lit2 migration conclusion, resume facsum

YZ

  • Previous:
    • Regular TMS and LBC work.
    • Some MODS and misc support stuff.
    • Investigating Python wrapper of the IIF interface (and python-irc program).
  • Next:
    • Mainly focus on TMS and LBC work.
    • Continue to work on some MODS utilities.
    • May refine the python-irc repo.
    • Other possible misc support stuff.

DE

  • Last week:
    • Autodesk licensing. Still not resolved for 2020, but >2020 works. Does not currently appear to be a showstopper.
      • Since I've not heard anything, I will contact the person who helped me with Vault licensing.
    • lucix disk swap
    • Started conversation with Glenn about hardware refresh
      • Everyone calls them "capture cards", but they're just fiber Ethernet to the Luci instrument network?
      • Assuming copper to the mountain network, how many fiber ports do we need in each machine?
    • SHARK + archive stuff, meeting set
    • Google SMTP relay issue, think it's UITS, but nothing from them so far. Will follow up
    • Various departures; waiting on management for offboarding go-ahead
  • This week
    • Aforementioned followups
    • JW and related fallout still pending
    • Refine Luci compute hardware

GE

  • Last week:
    • Discovered that lucix had a bad disk drive (/dataA). LeRoy and David E. replaced the drive.
      • Also discovered that lucix had been trying to warn us for a couple of years, and directed the lucifer and root mail to myself until I clean up any cron job spam; then will direct it to italerts+luci.
    • Fixed the FILENAME FITS header issue in IT 8684 by saving CUBE images to a subdirectory on luci.luci (but not in the archive, of course!).
    • More discussion of 8684 (EXPTIME FITS header), but no changes are needed.
    • Created a new test of LUCI double-buffering after eliminating redundant GEIRS status and other commands.
  • This week:
    • I encourage everyone to read the People's CDC Weather Reports.
    • Continue testing the fix for 8684 and GEIRS double-buffering.
    • Upgrade the osa-tcs TCS simulator to the latest TCS version for SHARK-NIR.
    • IT 8571, 8529 (Observer Scheduler GUI).

SCPH

  • SCADA system wth Dan R.
  • Core switches
  • lists for ITGC
  • HBPLC

next:

  • SCADA system
  • more lists for ITGC
  • mountain trip
    • core switch - check and install
    • obs rearrangement
    • particle monitor
    • printer
Topic revision: r9 - 08 Aug 2022, GlennEychaner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback